OpenStudy (anonymous):

Use substitution to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) y = 3x + 5 x = -3y - 5

OpenStudy (sleepyjess):

First, let's solve for y by substituting x into the equation. y = 3x + 5 y = 3(-3y-5) + 5 y = -9y - 15 + 5 y = -9y -10 y + 9y = -9y + 9y -10 10y = -10 10y/10 = -10/10 y = -1 Next, we solve for x by substituting y into the equation x = -3y - 5 x = -3(-1) - 5 x = 3-5 x = -2 to double check the work, plug x into the first equation y = 3x + 5 y = 3(-2) + 5 y = -6 + 5 y = -1 which is what we came up with above so you have x = -2 and y = -1 (-2, -1)

OpenStudy (sleepyjess):

You could have also solved for x first and done the same thing